In an experiment, 0.520 g of chromium metal reacts with 3.807 g of iodine. Using the conservation of mass law, predict the mass of product.


2 Cr(s) + 3 I2(s) 2 CrI3(s)
1
Expert's answer
2012-09-14T09:03:25-0400
n of Cr 0.52/52 = 0.01 mol
n I2 3.807/(127*2) = 0.015 mol

0.01 0.015 x
2Cr(s) + 3I2(s) --> 2CrI3(s)
2 3 2

x = 0.01 mol

m = 0.01* (52+3*127) = 4.33 g